# Self-gravitating solutions in Yang-Mills–Chern-Simons theory coupled to 3D massive gravity

### v1

@@ Abstract
We study self-gravitating solutions of 3-dimensional massive gravity coupled to Yang-Mills-Chern-Simons gauge theory. Among these, there is a family of asymptotically Warped-Anti de Sitter black holes that come to generalize previous solutions found in the literature and studied in the context of WAdS3/CFT2. We also present self-gravitating solutions to the 3-dimensional Einstein-Yang-Mills theory, as well as other self-gravitating solutions in the presence of higher-curvature terms.
@@ I Introduction
The 3-dimensional gravity provides a tractable scenario to explore many aspects of gravity that are of insurmountable technical complexity in four and higher dimensions. Well known examples of this are the microscopic derivation of the black hole entropy beyond supersymmetry Strominger (1998) and the explicit computation of the gravity partition function Maloney and Witten (2010). At classical level, the 3-dimensional models provide a useful playground to study the non-linear regime of gravity and gauge theories; in particular, in the context of black hole physics. This is mainly due to the existence of the Bañados-Teitelboim-Zanelli (BTZ) solution Bañados et al. (1992), which describes black holes with physical properties that mimic those of their higher-dimensional analogs but, at the same time, resulting much more tractable. Being locally equivalent to 3-dimensional Anti-de Sitter (AdS3) space Bañados et al. (1993), the BTZ geometry also appears in any other theory that admits an AdS3 vacuum. This includes 3-dimensional massive gravity, higher-spin theories, string theory and many others. A particularly interesting model to work with is the parity-even theory of massive gravity introduced in reference Bergshoeff et al. (2009a), which is commonly known as New Massive Gravity (NMG). This is a higher-curvature theory that propagates two massive, spin-2 local degrees of freedom and exhibits interesting properties, such as the existence of a rich variety of black holes, apart from the BTZ solution. This includes generalizations of the BTZ black holes with a weakened AdS3 asymptotics Oliva et al. (2009); Bergshoeff et al. (2009b), black holes with anisotropic scale invariance at infinity Ayón-Beato et al. (2009a), black holes in de Sitter (dS) space Oliva et al. (2009), and black holes in Warped AdS3 (WAdS3) spaces Clément (2009). All these solutions are of interest in the context of holography, as they were used to explore to what extent the AdS/CFT duality could be generalized beyond asymptotically AdS3 spaces Giribet et al. (2009); de Buyl et al. (2014); Gonzalez et al. (2011); Donnay and Giribet (2015, 2017); Chernicoff et al. (2024). Here, we will focus on the WAdS3 spaces.
@@
The WAdS3 spaces appear in connection to black hole physics: they emerge in the near-horizon region of Kerr black holes Bengtsson and Sandin (2006); Bardeen and Horowitz (1999) and are crucial ingredients in the formulation of the Kerr/CFT correspondence Guica et al. (2009). Besides, the WAdS3 spaces are at the root of the investigation of an entirely new class of conformal field theories (CFT) known as Warped-CFTs (WCFT) Detournay et al. (2012); see also Donnay and Giribet (2015, 2017). Consequently, any generalization of WAdS3 solutions, e.g. by coupling gauge fields, is in principle of interest both for black holes physics and for holography. Here, we will show that, when coupled to non-Abelian gauge theory, NMG admits black holes that asymptote WAdS3 spaces. These are similar to the WAdS3 black holes studied in Moussa et al. (2008) and Clément (2009), but, in contrast to those, the ones we construct here support a non-Abelian gauge field configuration. In fact, we will show that the gauge field may suffice to support the WAdS3 black holes even in absence of higher-curvature terms. The model we will consider consists of Yang-Mills–Chern-Simons (YM-CS) gauge theory coupled to NMG. This results in a higher-derivative theory of gravity with a non-linear source, and that is the reason why, in order to solve the problem, we need a good strategy.
@@
This strategy sounds simple; however, when trying to carry this out, one encounters a difficulty: Although WAdS3 spaces are solvable, they are not Einstein manifolds, i.e. they are not of constant curvature, and they are not even conformally flat. In addition, they are not compact and of negative curvature. All this implies that, while not intractable, the metric of WAdS3 black holes is much more involved than that of stationary BTZ. Besides, the theory we are dealing with is of fourth order in the metric, and non-linear both in the gravitational and in the matter sector. Thus, in order to circumvent the intrinsic difficulty of the problem, we find it convenient to look for these solutions progressively: In section II, we introduce the theory. In section III, we look for YM-CS solution on the 3-sphere with non-trivial gauge field configuration for the $SU(2)$ gauge field. The idea is to gain intuition from the compact case to select the right ansatz. In order to find such solutions, we propose a particularly friendly meron-type ansatz de Alfaro et al. (1976) for the gauge field, which is shown to suffice. Then, we appropriately introduce a real parameter and obtain a squashed 3-sphere solution in the theory. Finally, in section IV, after having gained some intuition about the squashing procedure, we perform the analytic continuation to the non-compact $SU(1,1)$ group, which amounts to work out the WAdS3 case, which has negative curvature. In this way, we finally find the WAdS3 black holes that we are interested in. It turns out that the meron-type ansatz results in a useful trick to find exact solutions of YM-CS theory coupled to gravity. In section V, which we reserve for some final remarks, we emphasize this point by presenting more new solutions of NMG coupled to YM-CS gauge theory.
@@ II Massive Gravity coupled to gauge theory
where $\kappa=(16\pi G)^{-1}$ is the gravitational coupling, $\Lambda$ is the cosmological constant, and $\mu$ is a real, dimension-1 parameter that represents the graviton mass. In the matter sector, $e$ denotes the coupling constant of the gauge theory. The non-Abelian field strength is defined as $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}+[A_{\mu},A_{\nu}]$, which, expressed in terms of differential forms, yields the curvature 2-form $F=\tfrac{1}{2}F_{\mu\nu}\text{d}x^{\mu}\wedge\text{d}x^{\nu}=\text{d}A+\tfrac{1}{2}[A,A]$. The gauge field decomposes as usual, $A=A_{\,\mu}\text{d}x^{\mu}=A^{i}_{\,\mu}t_{i}\text{d}x^{\mu}$, with $\mu=0,1,2$ being the spacetime indices and $i=1,2,3$ the group indices; $t_{i}$ are the generators of ${su}(2)$ algebra. The action also includes a CS term
@@
$\mathcal{E}^{(g)}_{\mu\nu}$ $\equiv R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R+\Lambda g_{\mu\nu}-\frac{1}{2\mu}K_{\mu\nu}-\frac{1}{2\kappa}T_{\mu\nu}=0\,,$ (3a)
@@ III Self-gravitating solutions with $SU(2)$ gauge field
As explained in the introduction, our first step will be finding solutions of positive curvature with non-trivial $SU(2)$ field configuration. That is to say, we look for solutions to Eqs. (3) by considering a ${su}(2)$-valued gauge connection $A=A_{\mu}^{i}t_{i}\text{d}x^{\mu}$, where $i=1,2,3$ are indices of the group, $t_{i}=-\frac{i}{2}\tau_{i}$ are generators of the $SU(2)$ satisfying the ${su}(2)$ Lie algebra $[t_{i},t_{j}]=\epsilon_{ijk}t^{k}$, and $\tau_{i}$ are the Pauli matrices. Notice that we have chosen the generators to be anti-hermitian. The standard (anti)commutation relations of the latter yield $\operatorname{Tr}(t_{i}t_{j})=-\tfrac{1}{2}\delta_{ij}$ and $\operatorname{Tr}(t_{i}t_{j}t_{k})=-\frac{1}{4}\epsilon_{ijk}$.
@@ III.1 The round 3-spheres
$\text{d}s^{2}=\rho^{2}\;\delta_{ij}\,\sigma^{i}\otimes\sigma^{j}=\rho^{2}\left[\left(\text{d}\psi+\cos\vartheta\text{d}\varphi\right)^{2}+\text{d}\theta^{2}+\sin^{2}\vartheta\text{d}\varphi^{2}\right]\,,$ (7)
where $\rho\in\mathbb{R}_{+}$ is a constant that denotes the radius of ${S}^{3}$. The metric is invariant under the action of the $SO(4)$ isometry group. For the non-Abelian field $A$, we assume an ansatz aligned along the left-invariant forms, i.e.
@@
$\lambda$ $=\frac{1}{2}+k\rho\,,$ (10a)
$\rho^{2}_{\pm}$ $=\frac{1}{4\left(4\kappa\Lambda e^{2}-k^{4}\right)}\left\{2\kappa e^{2}-k^{2}\pm e\sqrt{\kappa\left[4\left(\Lambda+\kappa e^{2}-k^{2}\right)-\frac{1}{\mu}\left(4\kappa\Lambda e^{2}-k^{4}\right)\right]}\right\}\,,$ (10b)
for $4\kappa\Lambda e^{2}-k^{4}\neq 0$. Notice that, in dimension 3, the presence of the CS term produces a shift in the value $\lambda=1/2$. This was first observed in the context of the 3-dimensional Einstein-YM-CS theory in Ref. Canfora et al. (2017). In fact, our solution (10) reduces to those found in Canfora et al. (2017) in the limit $\mu\to\infty$. Notice that, for $\rho_{\pm}^{2}$ to be positive, the parameters of the theory have to be constrained. Moreover, in the topological limit $e\to 0$, $k\to 0$ with $|k/e|<0$, where the matter content reduces to CS, one recovers the pure gauge solutions $\lambda=0$ and $\lambda=1$. If the CS coupling $k$ vanishes, there exist a perturbative branch for which $\Lambda\to 0$, i.e.
$\lambda=\frac{1}{2}\,,\ \ \ \ \ \ \ \rho^{2}_{-}=\frac{1}{16\mu}-\frac{1}{16\kappa e^{2}}+\mathcal{O}(\Lambda)\,.$ (11)
@@
$\lambda=\frac{1}{2}+k\rho\,,\ \ \ \ \ \rho^{2}=\frac{\kappa e^{2}-\mu}{8\mu\left(2\kappa e^{2}-k^{2}\right)}=\frac{k^{4}-4\mu\Lambda}{16\mu k^{2}(k^{2}-2\Lambda)}\,.$ (12)
The gauge field configuration that supports this solution has a nontrivial topological charge, what can be seen by integrating the CS term (2) evaluated on the configuration (8). This yields
$I_{\rm CS}=\frac{1}{2}+\frac{3k\rho}{2}-2k^{3}\rho^{3}\,,$ (13)
which receives contributions from the CS term.
@@ III.2 The squashed 3-sphere
${\rm d}s^{2}=\rho^{2}\left(\sigma_{1}^{2}+\sigma^{2}_{2}+\alpha^{2}\sigma_{3}^{2}\right)=\rho^{2}\left[{\rm d}\theta^{2}+\sin^{2}\theta\,{\rm d}\phi^{2}+\alpha^{2}({\rm d}\psi+\cos\theta\,{\rm d}\phi)^{2}\right]\,.$ (14)
Here, $\rho$ and $\alpha$ denote the radius and the squashing parameter, respectively. This metric is no longer invariant under the isometry group $SO(4)=SU(2)\times SU(2)\,/\mathbb{Z}_{2}$; its symmetry group is broken down to $SU(2)\times U(1)$ by the introduction of the squashing; $\alpha$ can be seen as a symmetry-breaking parameter. The squashed 3-sphere is an $\eta$-Einstein space, which means that it is equipped with a contact 1-form, $\eta=\eta_{\mu}\text{d}x^{\mu}$, and it satisfies the condition Okumura (1962); Blair (2002)
@@
for some $a$ and $b$. These geometries have been studied in the context of NMG in Ref. Flores-Alfonso et al. (2021a). In the case of the squashed 3-sphere, the contact form is $\eta=\sigma_{3}$ and the values of $a$ and $b$ are
$a=\alpha^{2}(\alpha^{2}-1),\ \ \ \ \ b=\frac{2-\alpha^{2}}{2\rho^{2}}.$ (16)
@@
To obtain a squashed 3-sphere solution with non-trivial $SU(2)$ gauge field configuration in NMG, we might feel tempted to take a simple ansatz such as
$A$ $=2\lambda\,\Big{[}\left(\sin\phi\,{\rm d}\theta+\cos\phi\cos\theta\sin\theta\,{\rm d}\phi\right)t_{1}$
$\qquad-\left(\cos\phi\,{\rm d}\theta-\sin\phi\cos\theta\sin\theta\,{\rm d}\phi\right)t_{2}-\sin^{2}\theta\,{\rm d}\phi\,t_{3}\Big{]}\,.$ (17)
This choice, however, leads to a gauge field with a trivial topological charge, as the gauge connection depends only on two coordinates of the squashed 3-sphere. It is possible to obtain a solution with non-trivial topological charge by considering an anisotropic version of (8). This has been explored in Ref. Canfora and Corral (2023) and it is easy to prove that the same class of solutions works in NMG equations. However, for simplicity, here we will focus on the case of vanishing topological charge.
Inserting Eq. (14) and the ansatz (17) into the field equations (3), we find that the latter reduces to a system of algebraic equations that admit the following solution
$\lambda=\frac{1}{2}\;,\;\;\;\;\;\;\;\;\rho^{2}=\frac{\alpha^{2}}{4k^{2}}\,,$ (18)
with the coupling constants being related as follows
$\mu$ $=-\frac{k^{2}\kappa\,e^{2}\left(21\alpha^{4}-25\alpha^{2}+4\right)}{2\alpha^{4}e^{2}\kappa-2\alpha^{2}e^{2}\kappa+4k^{2}}\,.$ (20)
A real, Lorentzian version of the solution can also be obtained by performing the analytic continuation $\alpha\to i\alpha$.

### v2

# Self-gravitating solutions in Yang-Mills-Chern-Simons theory coupled to 3D massive gravity

@@ Abstract
We study self-gravitating solutions of 3-dimensional massive gravity coupled to the Yang-Mills-Chern-Simons gauge theory. Among these, there is a family of asymptotically Warped-Anti de Sitter black holes that come to generalize previous solutions found in the literature and studied in the context of WAdS3/CFT2. We also present self-gravitating solutions to 3-dimensional Einstein-Yang-Mills theory, as well other self-gravitating solutions in the presence of higher-curvature terms.
@@ I Introduction
Three-dimensional gravity provides a tractable scenario to explore many aspects of gravity that are of insurmountable technical complexity in four and higher dimensions. Well known examples of this are the microscopic derivation of the black hole entropy beyond supersymmetry Strominger (1998) and the explicit computation of the gravity partition function Maloney and Witten (2010). At a classical level, the 3-dimensional models provide a useful playground to study the non-linear regime of gravity and gauge theories; in particular, in the context of black hole physics. This is mainly due to the existence of the Bañados-Teitelboim-Zanelli (BTZ) solution Bañados et al. (1992), which describes black holes with physical properties that mimic those of their higher-dimensional analogs but, at the same time, resulting much more tractable. Being locally equivalent to 3-dimensional Anti-de Sitter (AdS3) space Bañados et al. (1993), the BTZ geometry also appears in any other theory that admits an AdS3 vacuum. This includes 3-dimensional massive gravity, higher-spin theories, string theory and many others. A particularly interesting model to work with is the parity-even theory of massive gravity introduced in reference Bergshoeff et al. (2009a), which is commonly known as New Massive Gravity (NMG). This is a higher-curvature theory that propagates two massive, spin-2 local degrees of freedom and exhibits interesting properties, such as the existence of a rich variety of black holes, apart from the BTZ solution. This includes generalizations of the BTZ black holes with a weakened AdS3 asymptotics Oliva et al. (2009); Bergshoeff et al. (2009b), black holes with anisotropic scale invariance at infinity Ayón-Beato et al. (2009a), black holes in de Sitter (dS) space Oliva et al. (2009), and black holes in Warped AdS3 (WAdS3) spaces Clément (2009). All these solutions are of interest in the context of holography, as they were used to explore to what extent the AdS/CFT duality could be generalized beyond asymptotically AdS3 spaces Giribet et al. (2009); de Buyl et al. (2014); Gonzalez et al. (2011); Donnay and Giribet (2015, 2017); Chernicoff et al. (2024). Here, we will focus on the WAdS3 spaces.
@@
The WAdS3 spaces appear in connection to black hole physics: they emerge in the near-horizon region of Kerr black holes Bengtsson and Sandin (2006); Bardeen and Horowitz (1999) and are crucial ingredients in the formulation of the Kerr/CFT correspondence Guica et al. (2009). Besides, the WAdS3 spaces are at the root of the investigation of an entirely new class of conformal field theories (CFT) known as Warped-CFTs (WCFT) Detournay et al. (2012); see also Donnay and Giribet (2015, 2017). Consequently, any generalization of WAdS3 solutions, e.g. by coupling gauge fields, is in principle of interest both for black hole physics and for holography. Here, we will show that, when coupled to non-Abelian gauge theory, NMG admits black holes that asymptote WAdS3 spaces. These are similar to the WAdS3 black holes studied in Moussa et al. (2008) and Clément (2009), but, in contrast to those, the ones we construct here support a non-Abelian gauge field configuration. In fact, we will show that the gauge field may suffice to support the WAdS3 black holes even in absence of higher-curvature terms. The model we will consider consists of Yang-Mills–Chern-Simons (YM-CS) gauge theory coupled to NMG. This results in a higher-derivative theory of gravity with a non-linear source, and that is the reason why, in order to solve the problem, we need a good strategy.
@@
This strategy sounds simple; however, when trying to carry this out, one encounters a difficulty: Although WAdS3 spaces are solvable, they are not Einstein manifolds, i.e. they are not of constant curvature, and they are not even conformally flat. In addition, they are not compact and of negative curvature. All this implies that, while not intractable, the metric of WAdS3 black holes is much more involved than that of stationary BTZ. Besides, the theory we are dealing with is of fourth order in the metric, and non-linear both in the gravitational and in the matter sector. Thus, in order to circumvent the intrinsic difficulty of the problem, we find it convenient to look for these solutions progressively: In section II, we introduce the theory. In section III, we look for YM-CS solution on the 3-sphere with non-trivial gauge field configuration for the $SU(2)$ gauge field. The idea is to gain intuition from the compact case to select the right ansatz. In order to find such solutions, we propose a particularly friendly meron-type ansatz de Alfaro et al. (1976) for the gauge field, which is shown to suffice. Then, we appropriately introduce a real parameter and obtain a squashed 3-sphere solution in the theory. The newly added degree of anisotropy requires us to modify the gauge potential accordingly. In turn, this change requires us to solve higher-order polynomials in order to present explicit solutions Canfora and Corral (2023). Nonetheless, turning off the Chern-Simons term circumvents that difficulty. Thus, we concentrate on that case to present analytic anisotropic solutions. Finally, in section IV, after having gained some intuition about the squashing procedure, we perform the analytic continuation to the non-compact $SU(1,1)$ group, which amounts to work out the WAdS3 case, which has negative curvature. In this way, we finally find the WAdS3 black holes that we are interested in. It turns out that the meron-type ansatz results in a useful trick to find exact solutions of YM-CS theory coupled to gravity. We emphasize this point by presenting more new solutions of NMG coupled to YM-CS gauge theory in the Appendix. We reserve section V for some final remarks.
@@ II Massive Gravity coupled to gauge theory
where $\kappa=(16\pi G)^{-1}$ is the gravitational coupling, $\Lambda$ is the cosmological constant, and $\mu$ is a dimension-1 parameter. We use the standard notation for this last coupling constant, $\mu^{2}$, which after linearization represents the graviton mass. In the matter sector, $e$ denotes the coupling constant of the gauge theory. The non-Abelian field strength is defined as $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}+[A_{\mu},A_{\nu}]$, which, expressed in terms of differential forms, yields the curvature 2-form $F=\tfrac{1}{2}F_{\mu\nu}\text{d}x^{\mu}\wedge\text{d}x^{\nu}=\text{d}A+\tfrac{1}{2}[A,A]$. The gauge field decomposes as usual, $A=A_{\,\mu}\text{d}x^{\mu}=A^{i}_{\,\mu}t_{i}\text{d}x^{\mu}$, with $i=1,2,3$ being the group indices; $t_{i}$ are the generators of ${su}(2)$ algebra. When considering ${su}(1,1)$ algebras, we use $\eta_{i}$ for the generators instead. The action also includes a CS term
@@
$\mathcal{E}^{(g)}_{\mu\nu}$ $\equiv R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R+\Lambda g_{\mu\nu}-\frac{1}{2\mu^{2}}K_{\mu\nu}-\frac{1}{2\kappa}T_{\mu\nu}=0\,,$ (3a)
@@ III Self-gravitating solutions with $SU(2)$ gauge field
As explained in the introduction, our first step will be finding solutions of positive curvature with non-trivial $SU(2)$ field configuration. That is to say, we look for solutions to Eqs. (3) by considering an ${su}(2)$-valued gauge connection $A=A_{\mu}^{i}t_{i}\text{d}x^{\mu}$, where $i=1,2,3$ are indices of the group, $t_{i}=-\frac{i}{2}\tau_{i}$ are generators of the $SU(2)$ group satisfying the ${su}(2)$ Lie algebra $[t_{i},t_{j}]=\epsilon_{ijk}t^{k}$, and $\tau_{i}$ are the Pauli matrices. Notice that we have chosen the generators to be anti-hermitian. The standard (anti)commutation relations of the latter yield $\operatorname{Tr}(t_{i}t_{j})=-\tfrac{1}{2}\delta_{ij}$ and $\operatorname{Tr}(t_{i}t_{j}t_{k})=-\frac{1}{4}\epsilon_{ijk}$.
@@ III.1 The round 3-sphere
$\text{d}s^{2}=\rho_{0}^{2}\;\delta_{ij}\,\sigma^{i}\otimes\sigma^{j}=\rho_{0}^{2}\left[\left(\text{d}\psi+\cos\vartheta\text{d}\varphi\right)^{2}+\text{d}\vartheta^{2}+\sin^{2}\vartheta\text{d}\varphi^{2}\right]\,,$ (7)
where $\rho_{0}\in\mathbb{R}_{+}$ is a constant that denotes the radius of ${S}^{3}$. The metric is invariant under the action of the $SO(4)$ isometry group. For the non-Abelian field $A$, we assume an ansatz aligned along the left-invariant forms, i.e.
@@
$\lambda$ $=\frac{1}{2}+k\rho_{0}\,,$ (10a)
$\rho^{2}_{0}$ $=\frac{1}{4\left(4\kappa\Lambda e^{2}-k^{4}\right)}\left\{2\kappa e^{2}-k^{2}\pm e\sqrt{\kappa\left[4\left(\Lambda+\kappa e^{2}-k^{2}\right)-\frac{1}{\mu^{2}}\left(4\kappa\Lambda e^{2}-k^{4}\right)\right]}\right\}\,,$ (10b)
for $4\kappa\Lambda e^{2}-k^{4}\neq 0$. Notice that, in dimension 3, the presence of the CS term produces a shift in the value $\lambda=1/2$. This was first observed in the context of the 3-dimensional Einstein-YM-CS theory in Ref. Canfora et al. (2017). In fact, our solution (10) reduces to those found in Canfora et al. (2017) in the limit $\mu\to\infty$. Notice that, for $\rho_{0}^{2}$ to be positive, the parameters of the theory have to be constrained. Moreover, in the topological limit $e\to 0$, $k\to 0$ with $|k/e|<\infty$, where the matter content reduces to CS, one recovers the pure gauge solutions $\lambda=0$ and $\lambda=1$. If the CS coupling $k$ vanishes, there exist a perturbative branch for which $\Lambda\to 0$, i.e.
$\lambda=\frac{1}{2}\,,\ \ \ \ \ \ \ \rho_{0}^{2}=\frac{1}{16\mu^{2}}-\frac{1}{16\kappa e^{2}}+\mathcal{O}(\Lambda)\,.$ (11)
@@
$\lambda=\frac{1}{2}+k\rho_{0}\,,\ \ \ \ \ \rho_{0}^{2}=\frac{\kappa e^{2}-\mu^{2}}{8\mu^{2}\left(2\kappa e^{2}-k^{2}\right)}=\frac{k^{4}-4\mu^{2}\Lambda}{16\mu^{2}k^{2}(k^{2}-2\Lambda)}\,.$ (12)
The gauge field configuration that supports this solution has a nontrivial topological charge, which can be seen by integrating the CS term (2) evaluated on the configuration (8), giving
$I_{\rm CS}=\frac{1}{2}+\frac{3k\rho_{0}}{2}-2k^{3}\rho_{0}^{3}\,.$ (13)
@@ III.2 The squashed 3-sphere
${\rm d}s^{2}=\rho_{0}^{2}\left(\sigma_{1}^{2}+\sigma^{2}_{2}+\alpha^{2}\sigma_{3}^{2}\right)=\rho_{0}^{2}\left[{\rm d}\vartheta^{2}+\sin^{2}\vartheta\,{\rm d}\varphi^{2}+\alpha^{2}({\rm d}\psi+\cos\vartheta\,{\rm d}\varphi)^{2}\right]\,,$ (14)
with $\rho_{0}$ being the radius of the squashed 3-sphere. This metric is no longer invariant under the action of the isometry group $SO(4)=SU(2)\times SU(2)\,/\mathbb{Z}_{2}$; its symmetry group is broken down to $SU(2)\times U(1)$ due to squashing; $\alpha$ can be seen as a symmetry-breaking parameter. The squashed 3-sphere is an $\eta$-Einstein space, which means that it is equipped with a contact 1-form, $\eta=\eta_{\mu}\text{d}x^{\mu}$, and it satisfies the condition Okumura (1962); Blair (2002)
@@
for some $a$ and $b$. These geometries have been studied in Refs. Flores-Alfonso et al. (2021a); Andrianopoli et al. (2023, 2024). In the case of the squashed 3-sphere, the contact form is $\eta=\sigma_{3}$ and the values of $a$ and $b$ are
$a=\alpha^{2}(\alpha^{2}-1),\ \ \ \ \ b=\frac{2-\alpha^{2}}{2\rho_{0}^{2}}.$ (16)
@@
To obtain a squashed 3-sphere solution with non-trivial $SU(2)$ gauge field configuration in NMG, we take
$A=\sum_{i=1}^{3}\lambda_{(i)}\sigma_{i}t_{i}\,,$ (17)
as has been explored in Einstein gravity Canfora and Corral (2023). There, the solution for the YM-CS equations was found and remains unchanged in the NMG case. The relevant point here is that the matter equations fix the parameter $\lambda_{(i)}$ in terms of the squashing parameter and the CS coupling. The latter enters into the gravity sector only through $\lambda_{(i)}$ since, recall, the CS term is absent from the energy-momentum tensor due to its topological nature.
Although analytic solutions to the YM-CS equations can be obtained for $\lambda_{(i)}$, the metric cannot be obtained in the same fashion as (10) due to the presence of fifth-order polynomials in $\rho_{0}$. However, for $k=0$, the result is simply
$\lambda_{(1)}^{2}$ $=\lambda_{(2)}^{2}=\frac{\alpha}{8}\left(\sqrt{9\alpha^{2}+16}-3\alpha\right)\,,$ (18a)
$\lambda_{(3)}$ $=1-2\lambda_{(2)}^{2}\,,$ (18b)
which has nontrivial topological charge as shown in Ref. Canfora and Corral (2023). Inserting Eq. (14) and the ansatz (17) into the field equations (3), we find that solutions to the NMG equations sourced by YM matter are given by
$\rho_{0}^{2}$ $=\frac{(\lambda_{(2)}-\lambda_{(2)}\lambda_{(3)})^{2}-\alpha^{2}(\lambda_{(2)}^{2}-\lambda_{(3)})^{2}}{2\kappa e^{2}\alpha^{2}(\alpha^{2}-1)}+\frac{4-21\alpha^{2}}{8\mu^{2}}\,,$ (19a)
$\Lambda$ $=\frac{\alpha^{2}}{4\rho_{0}^{2}}+\frac{63\alpha^{4}-80\alpha^{2}+16}{64\mu^{2}\rho_{0}^{4}}+\frac{(\lambda_{(2)}^{2}-\lambda_{(3)})^{2}}{4\kappa e^{2}\rho_{0}^{4}}\,.$ (19b)
These results showcase how anisotropic meron type gauge fields are able to source gravitational backgrounds with a similar kind of anisotropy. Hence, we expect similar fields to exist for WAdS3 spacetimes.